# Semantic-guided modeling of spatial relation and object co-occurrence for indoor scene recognition

### v4

@@ Abstract
Exploring the semantic context in scene images is essential for indoor scene recognition. However, due to the diverse intra-class spatial layouts and the coexisting inter-class objects, modeling contextual relationships to adapt various image characteristics is a great challenge. Existing contextual modeling methods for scene recognition exhibit two limitations: 1) They typically model only one type of spatial relationship (order or metric) among objects within scenes, with limited exploration of diverse spatial layouts. 2) They often overlook the differences in coexisting objects across different scenes, suppressing scene recognition performance. To overcome these limitations, we propose SpaCoNet, which simultaneously models Spatial relation and Co-occurrence of objects guided by semantic segmentation. Firstly, the Semantic Spatial Relation Module (SSRM) is constructed to model scene spatial features. With the help of semantic segmentation, this module decouples spatial information from the scene image and thoroughly explores all spatial relationships among objects in an end-to-end manner, thereby obtaining semantic-based spatial features. Secondly, both spatial features from the SSRM and deep features from the Image Feature Extraction Module are allocated to each object, so as to distinguish the coexisting object across different scenes. Finally, utilizing the discriminative features above, we design a Global-Local Dependency Module to explore the long-range co-occurrence among objects, and further generate a semantic-guided feature representation for indoor scene recognition. Experimental results on three widely used scene datasets demonstrate the effectiveness and generality of the proposed method.
@@ 1 Introduction
Scene recognition is a popular research topic in the field of computer vision, aiming to identify the scene categories presented in images, such as ”bedroom,” ”living room,” and ”kitchen.” It is a prerequisite for robots to perform tasks in unknown environments (Chen et al., 2018; Liu et al., 2022) and plays a key role in various fields such as autonomous driving and human-computer interaction (Khan et al., 2023; Xie et al., 2020). Indoor scene recognition presents greater challenges compared to its outdoor counterpart, primarily due to the intra-class diverse spatial layout and the coexisting objects in different scene categories, as illustrated in Fig. 1. Consequently, it is crucial to develop effective methods for indoor scene representation.
Fig. 1: Some examples of different scene datasets. Images in the bedroom and hospital room categories could easily be confused. This inter-class similarity, caused by coexisting objects across scenes, could also be found in the classroom and restaurant. In contrast, the simple composition makes the variation between the different outdoor scene categories quite obvious.
@@
Similar to our insight, several recent strategies have also focused on extracting intra-scene object information for enhancing scene recognition, yielding promising results. To model spatial context, (Song et al., 2019) uses an object-relation-object triplet to explore the spatial order relationship among objects, while (Chen et al., 2020; Zeng et al., 2022) assign object labels to backbone features and model the spatial metric relationship among different object-related features. Despite considerable advancements, these studies focus solely on one type of spatial relationship (i.e., order or metric). In reality, as illustrated in Fig. 2, spatial relationships encompass not only order and metric types but also topological relationships (Egenhofer, 1990). Relying on any single type of these relationships is insufficient for analyzing such diverse spatial relationships within scenes. Also, these approaches (Chen et al., 2020; Zeng et al., 2022) overlook that space-irrelevant information in backbone features (color, etc.) can also lead to loss degradation during training, which may impede optimization of the network’s ability to represent spatial contexts. Considering these limitations, this paper designs a Semantic Spatial Relation Module (SSRM) dedicated to representing spatial contexts. SSRM first decouples the spatial information from scenes through semantic segmentation, ensuring the purity of the network input, thereby thoroughly exploring all spatial relationships among objects in an end-to-end manner.
@@
In addition to spatial relationships, object co-occurrence across scenes is also a significant contributor to scene recognition. Several approaches (Cheng et al., 2018; Zhou et al., 2021; Song et al., 2019) conclude the discriminative objects linked to specific scene categories by computing object probability distribution within scenes. However, a significant challenge arises when discriminative objects are identical across different scenes (e.g., the bed in both the bedroom and the hospital room shown in Fig. 1). This paper proposes a novel idea to solve this challenge inspired by a notable human ability: even if the discriminative objects in the bedroom and the hospital room are identical, humans can still easily discern the differences between these two scenes because the same objects exhibit different characteristics in different scene classes (e.g., apparent variances between the beds in the above two scenes). Building on this observation, before exploring object co-occurrence, we design a Semantic Node Feature Aggregation Module to assign input scene-related features to objects, thus enabling the network to distinguish identical objects like humans. Subsequently, we construct a Global-Local Dependency Module to leverage attention mechanisms to establish long-range co-occurrence among object features, and further generate a semantic-guided feature representation for indoor scene recognition.
@@
    A Semantic Spatial Relation Module is designed to model diverse spatial layouts within scenes, which decouples the spatial information from the given image with the help of semantic segmentation, thus avoiding the negative effects of irrelevant information on spatial modeling. By using spatial information as input, this module adaptively explores all spatial relationships among objects in an end-to-end manner.
@@
A preliminary version of this work was presented in the conference paper (Song & Ma, 2023). We make significant extensions from various aspects as follows: 1) We introduce a richer type of contextual relationship, i.e., a long-range co-occurrence among objects, to provide more discriminative information for better scene recognition. 2) We optimize the use of spatial relationships among objects within scenes, and evaluate new datasets to demonstrate the generalization of the proposed approach. 3) We conduct comprehensive quantitative ablation studies to show the effectiveness of each new proposed module. 4) A series of qualitative visualization studies for the SpaCoNet are provided to present the characteristics of the proposed method.
@@ 2.2 Semantic segmentation-based modeling
Spatial context-based methods: Considering the large intra-class difference caused by the complex and variable spatial layout within scenes, some studies have explored the spatial relationships between objects for scene recognition. Specifically, Song et al. (Song et al., 2019) propose to explore the spatial order relationship among objects using the ”object-relationship-object” triplet, and they encode the triplet as a feature representation based on Recurrent Neural Network (RNN) to assist scene recognition. In addition, ARG-Net (Zeng et al., 2022) firstly detects the semantic information of regions within scenes through semantic segmentation; then takes the features of different semantic regions as graph nodes and the spatial metric relationship between these regions as graph edges; finally obtains discriminative representations based on the graph convolution for scene recognition. The above studies usually focus on only one spatial relation. Still, as illustrated in Fig. 2, the spatial relations within scenes are not limited to order and metric relations, but also involve topological relations (Egenhofer, 1990). Existing studies relying solely on one type of spatial relations are insufficient to fully explore the diverse spatial layouts within scenes. In contrast, this paper proposes a Semantic Spatial Relation Module, which takes spatial information directly as input and thoroughly explores all spatial relations among objects in an end-to-end manner, ensuring the full exploitation of spatial information within scenes.
@@ 3 Our method
In this section, we will first describe the motivation and overview of the proposed SpaCoNet, and then present the specific details of each module.
@@ 3.1 Overview
Scene recognition aims to assign predefined scene category labels to given images. This study focuses on indoor scenes captured from a human perspective using digital cameras or smartphones (Choe et al., 2021), such as bedrooms, kitchens, and bookstores, etc. As shown in Fig. 1, indoor scenes are complex, comprising structural elements (e.g., walls, floors, doors), furniture (e.g., sofas, tables, beds), and decorative items (e.g., lamps, pictures, plants). Overall, the intra-scene diverse spatial layout and the coexisting objects across scenes present significant challenges for indoor scene recognition. Therefore, as shown in Fig. 3, this paper proposes SpaCoNet, guided by semantic information, aiming to simultaneously model the Spatial layout and Co-occurrence of objects for indoor scene recognition.
Initially, we employ PlacesCNN (Zhou et al., 2017a) as the Image Feature Extraction Module (IFEM) to extract deep scene features, which can be regarded as the baseline. Next, considering the diverse spatial layouts, we construct a Semantic Spatial Relation Module (SSRM) dedicated to modeling spatial features within the scene. Previous studies (Chen et al., 2020; Zeng et al., 2022) tend to explore one type of spatial relation, which cannot fully explore the diverse spatial layouts within scenes. Unlike them, SSRM takes spatial information directly as input and thoroughly explores all spatial relations among objects in an end-to-end manner.
@@
Overall, our proposed SpaCoNe contains five modules: (a) Semantic Spatial Relation Module, (b) Image Feature Extraction Module, (c) Semantic Node Feature Aggregation Module, (d) Global-Local Dependency Module, and (e) Scene Recognition Module. (a) provides feature maps that characterize the spatial relationships among object regions within the scene. (b) provides feature maps from PlacesCNN, which is pretrained on the large dataset Places365 (Zhou et al., 2017a), to obtain the advanced representation of image information. The feature maps from (a) and (b), as well as the semantic segmentation label map of the input, are sent to (c). (c) performs feature aggregation on these two feature maps guided by the label map to generate two semantic feature sequences, which are sent to (d). (d) then explores the long-range co-occurrence among semantic feature sequence through the attention mechanism and modifies the global features with the obtained information. Finally, the modified features are fed into (e) to predict the scene category.
@@ 3.2 Semantic Spatial Relation Module
Spatial relationships, as discussed in (Egenhofer, 1990), encompass three types: topological relationships (e.g., a pillow being surrounded by a bed), order relationships (e.g., a chair is behind a table), and metric relationships (e.g., the distance between an artboard and a person). Relying solely on one type of relationship (Order (Song et al., 2019) or Metric (Zeng et al., 2022)) is insufficient for analyzing such diverse spatial layouts. Therefore, we construct the Semantic Spatial Relation module (SSRM) as a dedicated branch to model spatial relationships implicitly in an end-to-end manner.
@@
Fig. 4: Semantic Spatial Relation Module (SSRM). The part surrounded by the red dashed box represents the confidence filtering stage, which is used to address semantic ambiguities. The part surrounded by the purple dashed box represents the spatial context modeling stage, which is used to comprehensively extract spatial features from the provided spatial information.
@@ 4.5 Comparison with state-of-the-art methods
Furthermore, compared to methods (Cheng et al., 2018; Pal et al., 2019; López-Cifuentes et al., 2020; Zeng et al., 2022; Miao et al., 2021; Zhou et al., 2021; Chen et al., 2018) that also incorporate object information for scene recognition, SpaCoNet demonstrates enhanced performance, highlighting the effectiveness of simultaneously exploiting spatial relation and co-occurrence among objects. To be specific, methods (Chen et al., 2020; Zeng et al., 2022) perform scene recognition by modeling spatial metric relationships among object regions; however, spatial relationships within scenes are not limited to metric relationships. In contrast, SpaCoNet first decouples spatial information from the scene and then employs a specialized branch to thoroughly model spatial features. This process enables more effective adaptation to complex and variable spatial layouts, resulting in superior scene recognition. In addition, while methods (Pal et al., 2019; Zhou et al., 2021; Cheng et al., 2018) also assist scene recognition by modeling object co-occurrence, they overlook the challenges posed by discriminative objects that coexist across different scenes. In contrast, SpaCoNet assigns scene-related features to individual objects, thus distinguishing objects with identical semantics across scenes. Then, SpaCoNet models object co-occurrence through the attention mechanism, yielding more advanced recognition performance.
@@ 5 Conclusions
Initially, we construct the Semantic Spatial Relation Module (SSRM) to specifically model spatial contextual relationships within scenes. Visualization experiments demonstrate that using dedicated modules to adaptively explore spatial information end-to-end allows the network to focus on finer scene details. In SSRM, an Adaptive Confidence Filter (ACF) is designed to mitigate the impact of errors in semantic segmentation results. Ablation studies for SSRM confirm the ACF’s significant contribution to improving scene recognition performance. ACF offers a new solution for mitigating semantic ambiguity in semantic-guided methods.

### v5

@@ Abstract
Exploring the semantic context in scene images is essential for indoor scene recognition. However, due to the diverse intra-class spatial layouts and the coexisting inter-class objects, modeling contextual relationships to adapt various image characteristics is a great challenge. Existing contextual modeling methods for scene recognition exhibit two limitations: 1) They typically model only one type of spatial relationship (order or metric) among objects within scenes, with limited exploration of diverse spatial layouts. 2) They often overlook the differences in coexisting objects across different scenes, suppressing scene recognition performance. To overcome these limitations, we propose SpaCoNet, which simultaneously models Spatial relation and Co-occurrence of objects guided by semantic segmentation. Firstly, the Semantic Spatial Relation Module (SSRM) is constructed to model scene spatial features. With the help of semantic segmentation, this module decouples spatial information from the scene image and thoroughly explores all spatial relationships among objects in an implicit manner, thereby obtaining semantic-based spatial features. Secondly, both spatial features from the SSRM and deep features from the Image Feature Extraction Module are allocated to each object, so as to distinguish the coexisting object across different scenes. Finally, utilizing the discriminative features above, we design a Global-Local Dependency Module to explore the long-range co-occurrence among objects, and further generate a semantic-guided feature representation for indoor scene recognition. Experimental results on three widely used scene datasets demonstrate the effectiveness and generality of the proposed method.
@@ 1 Introduction
Scene recognition is a popular research topic in the field of computer vision, aiming to identify the scene categories presented in images, such as ”bedroom,” ”living room,” and ”kitchen.” It is a prerequisite for robots to perform tasks in unknown environments (Chen et al., 2018; Liu et al., 2022) and plays a key role in various fields such as autonomous driving and human-computer interaction (Khan et al., 2023; Xie et al., 2020). Indoor scene recognition presents greater challenges compared to its outdoor counterpart, primarily due to the intra-class [^1] In this paper, the term ”class” consistently refers to scene categories. diverse spatial layout and the coexisting objects in different scene categories, as illustrated in Fig. 1. Consequently, it is crucial to develop effective methods for indoor scene representation.
Fig. 1: Some examples of different scene datasets. Images in the bedroom and hospital room categories could easily be confused due to shared objects like beds. This inter-class similarity, caused by coexisting objects across scenes, could also be found in the classroom and restaurant (e.g., tables). In contrast, the simple composition makes the variation between the different outdoor scene categories quite obvious.
@@
Similar to our insight, several recent strategies have also focused on extracting intra-scene object information for enhancing scene recognition, yielding promising results. To model spatial context, (Song et al., 2019) uses an object-relation-object triplet to explore the spatial order relationship among objects, while (Chen et al., 2020; Zeng et al., 2022) assign object labels to backbone features and model the spatial metric relationship among different object-related features. Despite considerable advancements, these studies focus solely on one type of spatial relationship (i.e., order or metric). In reality, as illustrated in Fig. 2, spatial relationships encompass not only order and metric types but also topological relationships (Egenhofer, 1990). Relying on any single type of these relationships is insufficient for analyzing such diverse spatial relationships within scenes. Also, these approaches (Chen et al., 2020; Zeng et al., 2022) overlook that space-irrelevant information in backbone features (color, etc.) can also lead to loss degradation during training, which may impede optimization of the network’s ability to represent spatial contexts. Considering these limitations, this paper designs a Semantic Spatial Relation Module (SSRM) dedicated to representing spatial contexts. SSRM first decouples the spatial information from scenes through semantic segmentation, ensuring the purity of the network input, thereby thoroughly exploring all spatial relationships among objects implicitly.
@@
In addition to spatial relationships, object co-occurrence across scenes is also a significant contributor to scene recognition. Several approaches (Cheng et al., 2018; Zhou et al., 2021; Song et al., 2019) conclude the discriminative objects linked to specific scene categories through statistical analysis of object probability distributions within scenes. However, a significant challenge arises when discriminative objects are identical across different scenes (e.g., the bed in both the bedroom and the hospital room shown in Fig. 1). This paper proposes a novel idea to solve this challenge inspired by a notable human ability: even if the discriminative objects in the bedroom and the hospital room are identical, humans can still easily discern the differences between these two scenes because the same objects exhibit different characteristics in different scene classes (e.g., apparent variances between the beds in the above two scenes). Building on this observation, before exploring object co-occurrence, we design a Semantic Node Feature Aggregation Module to assign input scene-related features to objects, thus enabling the network to distinguish identical objects like humans. Subsequently, we construct a Global-Local Dependency Module to leverage attention mechanisms to establish long-range co-occurrence among object features, and further generate a semantic-guided feature representation for indoor scene recognition.
@@
    A Semantic Spatial Relation Module is designed to model diverse spatial layouts within scenes, which decouples the spatial information from the given image with the help of semantic segmentation, thus avoiding the negative effects of irrelevant information on spatial modeling. By using spatial information as input, this module thoroughly explores all spatial relationships among objects in an implicit manner.
@@
A preliminary version of this work was presented in the conference paper (Song & Ma, 2023). We make significant extensions from various aspects: 1) We introduce a richer type of contextual relationship, i.e., a long-range co-occurrence among objects, to provide more discriminative information for better scene recognition. 2) We optimize the use of spatial relationships among objects within scenes, and evaluate new datasets to demonstrate the generalization of the proposed approach. 3) We conduct comprehensive quantitative ablation studies to show the effectiveness of each new proposed module. 4) A series of qualitative visualization studies for the SpaCoNet are provided to present the characteristics of the proposed method.
@@ 2.2 Semantic segmentation-based modeling
Spatial context-based methods: Considering the large intra-class difference caused by the complex and variable spatial layout within scenes, some studies have explored the spatial relationships between objects for scene recognition. Specifically, Song et al. (Song et al., 2019) propose to explore the spatial order relationship among objects using the ”object-relationship-object” triplet, and they encode the triplet as a feature representation based on Recurrent Neural Network (RNN) to assist scene recognition. In addition, ARG-Net (Zeng et al., 2022) firstly detects the semantic information of regions within scenes through semantic segmentation; then takes the features of different semantic regions as graph nodes and the spatial metric relationship between these regions as graph edges; finally obtains discriminative representations based on the graph convolution for scene recognition. The above studies usually focus on only one spatial relation. Still, as illustrated in Fig. 2, the spatial relations within scenes are not limited to order and metric relations, but also involve topological relations (Egenhofer, 1990). Existing studies relying solely on one type of spatial relations are insufficient to fully explore the diverse spatial layouts within scenes. In contrast, this paper proposes a Semantic Spatial Relation Module, which takes spatial information directly as input and implicitly models spatial features end-to-end, ensuring thoroughly explores all spatial relations among objects within scenes.
@@ 3 Our method
Scene recognition aims to assign predefined scene category labels to given images. Formally, let $I\in{\mathbb{R}^{w\times h\times 3}}$ represent a given image, where $w$ and $h$ denote its width and height, respectively. The set of scene categories to be classified is defined as $T=\left\{t_{1},t_{2},...,t_{\left|T\right|}\right\}$. The objective is to predict the category of the given image $I$ using the following formula:
$class(I)=\arg\mathop{\max}\limits_{t\in T}(\text{softmax}({\Psi(I)}))_{t}$ (1)
where $\Psi$ represents the designed pipeline for extracting representative information from images (e.g., the SpaCoNet proposed in this paper). In the following, we will first describe the motivation and overview of the proposed SpaCoNet, and then present the specific details of each module.
@@ 3.1 Overview
This study focuses on indoor scenes captured from a human perspective using digital cameras or smartphones, such as bedrooms, kitchens, and bookstores, etc. As shown in Fig. 1, indoor scenes are complex, comprising structural elements (e.g., walls, floors, doors), furniture (e.g., sofas, tables, beds), and decorative items (e.g., lamps, pictures, plants). Overall, the intra-scene diverse spatial layout and the coexisting objects across scenes present significant challenges for indoor scene recognition. To meet these challenges, as shown in Fig. 3, this paper proposes SpaCoNet, guided by semantic information, aiming to simultaneously model the Spatial layout and Co-occurrence of objects to extract representative information for indoor scene recognition.
Initially, we employ PlacesCNN (Zhou et al., 2017a) as the Image Feature Extraction Module (IFEM) to extract deep scene features, which can be regarded as the baseline. Next, considering the diverse spatial layouts, we construct a Semantic Spatial Relation Module (SSRM) dedicated to modeling spatial features within the scene. Previous studies (Chen et al., 2020; Zeng et al., 2022) tend to explore one type of spatial relation, which cannot fully explore the diverse spatial layouts within scenes. In contrast, SSRM takes spatial information directly as input and implicitly models spatial features end-to-end, ensuring thoroughly explores all spatial relations among objects.
@@
Overall, to simultaneously model the Spatial layout and Co-occurrence of objects, our proposed SpaCoNet contains five modules. Initially, to thoroughly explore diverse spatial relations within scenes, we design the (a) Semantic Spatial Relation Module, which implicitly models spatial features end-to-end. Concurrently, (b) Image Feature Extraction Module processes the original image to fully utilize scene information. Furthermore, to address the challenge of modeling object co-occurrence in scenes where representative objects may be identical, the feature maps from (a) and (b) are passed to (c) Semantic Node Feature Aggregation Module. (c) assign input scene-related features to each object node, thereby distinguishing objects with identical semantics in different scenes, and generating semantic feature sequences. These sequences are then forwarded to (d) Global-Local Dependency Module, which explores long-range co-occurrence within semantic feature sequences through the attention mechanism and modifies the global features with the obtained information. Finally, the modified features are fed into (e) Scene Recognition Module to predict the scene category.
@@ 3.2 Semantic Spatial Relation Module
Spatial relationships, as discussed in (Egenhofer, 1990), encompass three types: topological relationships (e.g., a pillow being surrounded by a bed), order relationships (e.g., a chair is behind a table), and metric relationships (e.g., the distance between an artboard and a person). Relying solely on one type of relationship (Order (Song et al., 2019) or Metric (Zeng et al., 2022)) is insufficient for analyzing such diverse spatial layouts. Therefore, we construct the Semantic Spatial Relation module (SSRM) as a dedicated branch to model spatial relationships in an implicit manner.
@@
Fig. 4: Semantic Spatial Relation Module (SSRM). The part surrounded by the red dashed box represents the confidence filtering stage, which is used to address semantic ambiguities. The part surrounded by the purple dashed box represents the spatial context modeling stage, which is used to implicitly model spatial features from the provided spatial information.
@@ 4.5 Comparison with state-of-the-art methods
Furthermore, compared to methods (Cheng et al., 2018; Pal et al., 2019; López-Cifuentes et al., 2020; Zeng et al., 2022; Miao et al., 2021; Zhou et al., 2021; Chen et al., 2018) that also incorporate object information for scene recognition, SpaCoNet demonstrates enhanced performance, highlighting the effectiveness of simultaneously exploiting spatial relation and co-occurrence among objects. To be specific, methods (Chen et al., 2020; Zeng et al., 2022) perform scene recognition by modeling spatial metric relationships among object regions; however, spatial relationships within scenes are not limited to metric relationships. In contrast, SpaCoNet first decouples spatial information from the scene and then employs a specialized branch to thoroughly model spatial features. This process enables more effective adaptation to complex and variable spatial layouts, resulting in superior scene recognition. In addition, while some statistical methods (Pal et al., 2019; Zhou et al., 2021; Cheng et al., 2018) also assist scene recognition by modeling object co-occurrence, they overlook the challenges posed by discriminative objects that coexist across different scenes. In contrast, SpaCoNet assigns scene-related features to individual objects, thus distinguishing objects with identical semantics across scenes. Then, SpaCoNet models object co-occurrence through the attention mechanism, yielding more advanced recognition performance.
@@ 5 Conclusions
Initially, we construct the Semantic Spatial Relation Module (SSRM) to specifically model spatial contextual relationships within scenes. Visualization experiments demonstrate that using dedicated modules to implicitly model spatial information allows the network to focus on finer scene details. In SSRM, an Adaptive Confidence Filter (ACF) is designed to mitigate the impact of errors in semantic segmentation results. Ablation studies for SSRM confirm the ACF’s significant contribution to improving scene recognition performance. ACF offers a new solution for mitigating semantic ambiguity in semantic-guided methods.